# Robust Optimal Safe and Stability Guaranteeing Reinforcement Learning Control for Quadcopter

### v1

@@ III-B Lipschitz Continuity of NN Controller and its QC
where $L$ is the Lipschitz bound constant. The class of all control maps $\pi(\cdot)$ that are $L$-Lipschitz is denoted by $\Pi_{L}$.
Neural networks are composed of linear operations and nonlinear activation functions. The linear operations have bounded gradients, with their Lipschitz constant determined by the largest singular value of the weight matrix. On the other hand, the nonlinear activation functions such as, Sigmoid, ReLU, and tanh, have Lipschitz constants between 0 and 1 [13]. Computing the exact Lipschitz constant of an NN is challenging due to the piecewise application of activation functions, making the gradient calculation complex. However, efficient estimation of the Lipschitz constant for NNs can be achieved using semidefinite programming (SDP) solvers [5].
As shown in [1, Proposition 1], each output of an NN can be expressed as a linear combination of scalar functions and input elements, with partial derivatives bounded between $-L$ and $L$. As shown next, this result remains valid even when using the L2-norm, ensuring that the Lipschitz continuity of the NN remains appropriately represented in the analysis. From the definition of Lipschitz continuity (42) and triangular inequality, the NN controller $\pi_{NN}$ satisfies the following:
@@
$=L\sqrt{\sum_{j=1}^{n}{\lvert s_{1,j}-s_{2,j}\rvert}^{2}}$
$\leq L\sum_{j=1}^{n}|s_{1,j}-s_{2,j}|,\forall s_{1},s_{2}\in\mathbb{R}^{n}$
Also, the infinity-norm is always less than the 2-norm:
$\left\lVert\pi_{NN}(s_{1})-\pi_{NN}(s_{2})\right\rVert_{\infty}\leq\left\lVert\pi_{NN}(s_{1})-\pi_{NN}(s_{2})\right\rVert_{2}.$
From above two inequalities, the same result with [1, Proposition 1] is obtained even when we use 2-norm to define Lipschitz continuity (as opposed to the $\infty$-norm):
@@
It then follows that there exists a set of scalar functions: $\delta_{ij}:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow[-L,L],\forall i\in\{1,...,m\},\forall\in\{1,...,n\}$ such that $\forall s_{1},s_{2}\in\mathbb{R}^{n}$:
@@
which is the same core result in the proof of [1, Proposition 1]. Therefore, the QC of [1, Proposition 1] is still valid even under the use of L2-norm based Lipschitz continuity. Note the change in norm, results in a new version of the input domain implied by the state domain $\mathcal{S}$:
@@
III-C Sector bound of nonlinear and parameter variation and its QC
The system of (41) under nominal state-feedback control $\pi_{0}(s)=Ks$ can be decomposed into linear and “nonlinear and parameter variation (NPV)” components, by linearization of the nominal system around the equilibrium and considering the rest of the nonlinearity:
$\dot{s}(t)=\underbrace{(A_{0}+B_{0}K)}_{A_{K}}s(t)+\underbrace{B_{0}u_{NN}(t)+\eta(s(t),u_{NN}(t),\alpha(t))}_{{\rm NPV:}\eta(s(t),u_{NN}(t),\alpha(t))}$
where $u_{NN}(t)=\pi_{NN}(s(t))$, the entries in the linearized nominal part are obtained from the Jacobians of $f$ and $\pi$ (that exist owing to their local continuous differentiability assumption) as follows:
@@ III-C Sector bound of nonlinear and parameter variation and its QC
and $\eta(s,u_{NN},\alpha)=f(s,Ks+u_{NN},\alpha)-A_{K}s-B_{0}u_{NN}$ is the remaining additive nonlinear part. The overall NPV part is the component not included in the linearized nominal part, namely:
$\zeta(s(t),u_{NN}(t),\alpha(t))=B_{0}u_{NN}(t)+\eta(s(t),u_{NN}(t),\alpha(t)).$
@@
III-D LMI for robust stability
Using the QCs for the NN controller Lipschitz bound and system NPV sector bound, in [1], an LMI constraint is obtained to guarantee the existence of a robust Lyapunov function over a given state domain, Lipschitz constant, and parameter variation bound. The feasibility of this LMI constraint ensures the existence of a robust quadratic Lyapunov function to certify the the stability of the NN-controlled system and a RoS within the specified safety set that is invariant and ensures convergence to equilibrium, even under parameter variations.
@@ III-D LMI for robust stability
For the given system of (41), when the below LMI constraint of (51) is feasible for some $L,K,(P>0),(\Lambda\geq 0),(\gamma_{ij}\geq 0,\forall i\in 1,\ldots,m,j\in 1,\ldots,n)$ over the given state parameter variation domains, then there exists a robust quadratic Lyapunov function $\mathcal{V}(s)=s^{T}Ps$ such that $\pi(s)=Ks+\pi_{NN}(s);\,\pi_{NN}\in\Pi_{L}$ stabilizes the system of (41) over the given state domain under the variations of the parameters in its given domain.
@@
IV Robust Stabilizing Control for quadcopter
IV-A Dynamics with parameter variation & Nominal control
@@
TABLE I: Constant for Quadcopter Model
@@
A quadcopter is an underactuated system where the degree of freedom for control is less than the number of states. Given a desired position ($x_{d},y_{d},z_{d}$) and heading ($\psi_{d}$), a cascade proportional-derivative (PD) controller is employed as the nominal controller in our design. The positional P-controller computes the desired thrust ($T_{d}$) and roll/pitch angles ($\phi_{d},\theta_{d}$), which are then used by the attitude PD-controller to calculate the desired torque ($\tau_{d}$). The structure of this nominal controller is shown in Fig 2.
Figure 2: Nominal Controller
The control inputs are set as the motor thrusts, which serve as the actuator commands to stabilize and guide the quadcopter:
@@
The thrust and torque generated by the quadcopter’s rotors are proportional to the superposition of the square of rotor speeds. However, these values are influenced by aerodynamic conditions and external disturbances. To account for these uncertainties, a parameter variation is introduced into the thrust and torque model and compensated through the added NN controller. Table II represents the parameter variation bounds, indicating the extent to which actual thrust and torque may deviate from their commanded values. Modeling these variations ensures that the model accounts for the real-world discrepancies, enhancing robustness of the control design.
TABLE II: Parameter Variation
@@
The system equation (92) consists of 12 states, 4 control inputs, and 4 parameter variations. The coordinates are shifted such that the equilibrium point is at the origin, i.e., $(x_{d},y_{d},z_{d},\psi_{d})=(0,0,0,0)$. With this coordinate shift, so that $(x_{d},y_{d},z_{d},\psi_{d})=(0,0,0,0)$, it follows that nominal Position and Attitude controllers are state-feedback. Table III lists the parameter values for the nominal controller, which is chosen to ensure that the state matrix associated with the nominal controller, $A_{K}$, is stable with all its eigenvalues having negative real parts, i.e., the nominal controller stabilizes the linearized nominal system under zero parameter variation around the equilibrium point.
TABLE III: PD control gain
@@
IV-B Maximal Lipschitz Bound and Safety Domain
The feasibility of the LMI constraint (51) depends on system nonlinearity, safety domain, and parametric bound, and when feasible, one can find $(L,K,P,\Lambda,\{\lambda_{ij}\})$ that satisfy the LMI constraint so that the system (41) under the control $\pi(s)=Ks+\pi_{NN}(s);\,\pi_{NN}\in\Pi_{L}$ is certified as asymptotically stable with the Lyapunov function $\mathcal{V}=s^{T}Ps$ having a RS-RoS (Robust Safe Region of Stability) that is a sublevel set of $\mathcal{V}$. To maximize the class of robust, safe, and stable NN controllers, $L$ should be as large as possible, and to maximize the size of RS-RoS the corresponding safety domain $\mathcal{S}$ should also be as large as possible. Maximal ($L,\mathcal{S}$) values are determined through the iterative search method of [1, Algorithm 1]. The process begins with an initial guess ($L_{0},\mathcal{S}_{0}$) set to small values, which for our quadcopter case is set as below:
$L_{0}=1.0;\mathcal{S}_{0}\!\equiv\!\left[\!\!\begin{array}[]{c}\mathcal{S}_{0}(x)=\mathcal{S}_{0}(y)\\ \mathcal{S}_{0}(z)\\ \mathcal{S}_{0}(\phi)=\mathcal{S}_{0}(\theta)=\mathcal{S}_{0}(\psi)\\ \mathcal{S}_{0}(\dot{x})=\mathcal{S}_{0}(\dot{y})\\ \mathcal{S}_{0}(\dot{z})\\ \mathcal{S}_{0}(p)=\mathcal{S}_{0}(q)=\mathcal{S}_{0}(r)\end{array}\!\!\right]\!=\!\left[\!\!\begin{array}[]{c}(-1.5,1.5)\\ (-0.3,0.3)\\ (-0.2,0.2)\\ (-1.5,1.5)\\ (-0.3,0.3)\\ (-0.2,0.2)\end{array}\!\!\right]\!.$
The ($L,\mathcal{S}$) values are iteratively increased, and the feasibility of the LMI constraint is evaluated at each step. If the LMI constraint becomes infeasible, ($L,\mathcal{S}$) is reduced, and the feasibility is rechecked. This process is repeated with progressively smaller step sizes, until the algorithm converges to the maximal Lipschitz bound and safety domain, denoted as ($L^{*},\mathcal{S}^{*}$), which in our case is found to be:
$L^{*}\!=\!1.330;\mathcal{S}^{*}\!\!\equiv\!\!\!\left[\!\!\!\!\begin{array}[]{c}\mathcal{S}^{*}(x)\!=\!\mathcal{S}^{*}(y)\\ \mathcal{S}^{*}(z)\\ \mathcal{S}^{*}(\phi)\!=\!\mathcal{S}^{*}(\theta)\!=\!\mathcal{S}^{*}(\psi)\\ \mathcal{S}^{*}(\dot{x})\!=\!\mathcal{S}^{*}(\dot{y})\\ \mathcal{S}^{*}(\dot{z})\\ \mathcal{S}^{*}(p)\!=\!\mathcal{S}^{*}(q)\!=\!\mathcal{S}^{*}(r)\end{array}\!\!\!\!\right]\!\!\!=\!\!\!\left[\!\!\!\!\begin{array}[]{c}(-1.9951,1.9951)\\ (-0.3990,0.3990)\\ (-0.2660,0.2660)\\ (-1.9951,1.9951)\\ (-0.3990,0.3990)\\ (-0.2660,0.2660)\end{array}\!\!\!\!\right]\!\!.$
@@ IV-C Optimal NN Controller Training
To optimally train the NN controller satisfying the maximal Lipschitz bound $L^{*}$ found above to minimize control cost and reference tracking error, we employ Proximity Policy Optimization (PPO) [15], a deep reinforcement learning (RL) method based on an actor-critic structure. In an RL, an agent interacts with the environment to collect samples, where each sample is represented as $(s,a,r,\gamma,s^{\prime})$ where $s\in\mathcal{S},a\in\mathcal{A}$ are the current state and action ($\mathcal{S}$ and $\mathcal{A}$ are state and action domains), $r\in\mathbb{R}$ is the associated reward, $\gamma\in[0,1)$ is a discount factor, and $s^{\prime}\in\mathcal{S}$ is the next state. These samples represent the transition from $s$ to $s^{\prime}$ when the action $a$ taken, and the reward $r$ is received. Using the collected samples, policy and value networks are trained, where neural networks are used as function approximators, representing maps from state to action (actor network) and from state to value (value network), value is the accumulated total rewards (long-term expected return). (In contrast, the reward is an immediate return of a single step.) The goal of RL is to optimize the policy parameters to maximize value, thereby improving the agent’s performance in achieving its objectives.
Fig 3 illustrates this actor-critic framework used in the optimal NN training process. The environment models the quadcopter dynamics along with the nominal controller described in Equation (92). The NN controller provides control inputs that are fed into the environment, while the environment returns a reward value for each action depending on the current state, and also advances the state. The “advantage” is defined as the difference between the state value and an action value, providing a measure of whether a particular action is better or worse than the default action suggested by the policy. We estimated the advantage using the method proposed in [16], and computed the policy gradient based on this estimated advantage.
Figure 3: Actor Critic Framework
The reward indicates how effectively the agent performs in achieving the desired objective, guiding the optimization of the policy. We design a reward function to minimize the error between the reference state and the current state. When controlling the quadcopter, two aspects of performance are considered: 1) transient response and 2) steady-state response. When the quadcopter is far from the reference position, its reward is computed from only the positional and directional errors. Once the quadcopter is close to the reference position, angular velocity, that must become small, is also penalized (along with the positional and directional errors) to reduce the oscillations around the reference. Given the position and direction errors defined as $p_{e}:=[x_{d}-x,y_{d}-y,z_{d}-z]^{T}$ and $\psi_{e}:=\psi_{d}-\psi$, and the angular velocity defined by $\nu:=[p,q,r]^{T}$, the reward function is formulated to capture these objectives:
@@
Fig 4 shows the actor and critic network structures. The actor consists of five hidden layers, each containing 64 neurons with Rectified Linear Unit (ReLU) as the activation function. The final layer performs only linear operations without an activation function. The actor has two output branches: one computes the mean, and the other computes the variance of the action. Actions are sampled from a multivariate normal distribution using the mean and variance. During testing, the variance output is ignored, and deterministic actions (mean values) are used to control the quadcopter. The critic network consists of four hidden layers, each with 192 or 256 neurons and ReLU activation function. This network evaluates the value function, guiding the optimization of the actor network.
Figure 4: Actor and Critic networks
At the beginning of each episode, the quadcopter’s position is initialized at the origin, and the reference position is randomly selected within two-meter radius from the origin. Parameter variation ($\alpha$) is randomly sampled at every timestep from an uniform distribution within the given bounds, introducing uncertainties in thrust and torque. The simulation uses the first-order integration to update the state based on (92). Both actor and critic networks are trained using PPO. Table V provides the PPO parameters used for training. Definitions of these parameters can be found in [15], [16], and [17].
TABLE IV: PPO parameter
@@
To certify the asymptotic stability of the closed-loop system, the Lipschitz constant of the NN controller must remain below the maximal Lipschitz bound $L^{*}$ computed in Section IV-B. It is therefore important to obtain a tight estimation of an NN’s Lipschitz constant. To achieve this, we apply the semi-definite programming based estimation method proposed in [5] to estimate the Lipschitz constant of the trained actor network (without the variance branch). When the estimated Lipschitz bound of the trained NN is higher than the $L^{*}$ value, we scale down the weights of the final layer of the actor network so that the maximal Lipschitz bound is satisfied. In our study, the Lipschitz constant of the trained actor network came out to be $L=3.1519>L^{*}=1.3301$. So we scaled down the weight of the final layer by a factor smaller than $0.42200.4<L^{*}/L=$ 0.4220. After adjusting the weights of the NN controller, its Lipschitz constant is reduced to below the bound:
$L^{new}=1.2608.$
Note since in our design, the final layer is a linear operation without an activation function, scaling it down does not alter the action pattern inferred by the actor network. Instead, it simply reduces the magnitude of the outputs, ensuring that the Lipschitz constant remains below the required bound. This adjustment ensures the controller’s stability while preserving its functional behavior.
V Controller Evaluation
We evaluate the performance of the quadcopter under the nominal and NN controllers in simulation. To show the robustness of control, the parameter $\alpha$ is randomly drawn from the uniform distribution within the specified parameter bound.
V-A Reference Position Tracking
The reference position tracking capability of the quadcopter was evaluated under two scenarios: (1) using only the nominal controller, and (2) using both the nominal and NN controllers. The quadcopter was initialized at $p_{init}=[0,0,0]^{T}$ with zero initial attitude, velocity and angular velocity, while the reference position was set to $p_{d}=[1.4,0,0]^{T}$.
Control inputs from the nominal controller, the NN controller, and their combined outputs are shown in Figure 5. These inputs highlight the NN controller’s role in dynamically adjusting the system’s behavior to minimize position errors.

### v2

@@ III-B Lipschitz bounded control with New enhancements
where $L$ is termed the Lipschitz bound constant. The class of all control maps $\pi(\cdot)$ that are $L$-Lipschitz (Lipschitz bounded with bound $L$) is denoted by $\Pi_{L}$
As shown in [1, Proposition 1], each output of an $L$-Lipschitz NN can be expressed as a linear combination of scalar functions weighted by the input elements, with partial derivatives of the scalar functions bounded between ±$\pm L$. As shown next, this essential result remains valid even when using the L2-norm based definition of Lipschitz continuity, preserving all the ensuing results and the framework of [1]. From the definition of Lipschitz continuity (52), we have:
@@
$=L\sqrt{\sum_{j=1}^{n}{\lvert s_{1,j}-s_{2,j}\rvert}^{2}}\leq L\sum_{j=1}^{n}|s_{1,j}-s_{2,j}|,\forall s_{1},s_{2}\in\mathbb{R}^{n}$
where the last inequality follows from the triangular inequality. Also, since the infinity-norm is always less than the 2-norm:
$\left\lVert\pi_{NN}(s_{1})-\pi_{NN}(s_{2})\right\rVert_{\infty}\leq\left\lVert\pi_{NN}(s_{1})-\pi_{NN}(s_{2})\right\rVert_{2},$
we can conclude the same result as in the proof step of [1, Proposition 1] even with the use of L2-norm to define Lipschitz continuity (as opposed to the $\infty$-norm):
@@
It then follows that there exists scalar functions: $\delta_{ij}:\mathcal{S}\times\mathcal{S}\rightarrow[-L,L],\forall i\in\{1,...,m\},\forall\in\{1,...,n\}$ such that $\forall s_{1},s_{2}\in\mathcal{S}$:
@@
which again is the same core result in the proof of [1, Proposition 1]. Therefore, the QC of [1, Proposition 1] is still valid even under the use of L2-norm based Lipschitz continuity. Note the change in norm, results in a new version of the input domain implied by the state domain $\mathcal{S}$ that is used below to obtain a local sector bound for system nonlinearity and parameter variability:
@@
A second enhancement we introduce is the manner in which the Lipschitz bound of an NN is estimated. Neural networks are composed of linear operations of weighted sums and nonlinear activations. Under the L2-norm based definition of Lipschitz continuity utilized above, the Lipschitz constant for the linear operators is simply determined by the largest singular values of the corresponding weight matrices [20]. On the other hand, the nonlinear activation functions, such as Sigmoid, ReLU, and tanh, have Lipschitz constants between 0 and 1. Computing an exact Lipschitz constant of an NN is, in general, challenging due to the piecewise application of the activation functions, which makes the computation of the overall gradient bound complex. In fact, the work presented in [1] presented a rather conservative estimate, taken to be the product of the largest absolute row sum norm of weights of each layer [20]. In this work, we employ a semidefinite programming (SDP) based approach [5] for estimating the Lipschitz constant for NNs that is much less conservative and remains computationally efficient.
A third enhancement we introduce is the way to impose a desired Lipschitz bound for an NN controller: We keep the last layer linear, free from activation function, and simply scale down the weights of the last layer proportionately so as to achieve a desired Lipschiz bound, without affecting the weights of the inner layers of an otherwise optimal NN. The results for a real-life example of a quadcopter control are presented in Section IV.
III-C Sector bound of nonlinearity and parameter variation
The system of (51) under state-feedback control $u(s)=\pi(s)=\pi_{0}(s)+\pi_{NN}(s)=Ks+u_{NN}(s)$ can be decomposed into linear and nonlinear and parameter variation (NPV) components, by linearization of the nominal system around the equilibrium and considering the rest of the nonlinearity:
@@ III-C Sector bound of nonlinearity and parameter variation
$=\underbrace{(A_{0}+B_{0}K)}_{A_{K}}s+\underbrace{f(s,u(s),\alpha)-(A_{0}+B_{0}K)s}_{{\rm NPV:}\zeta(s,u_{NN}(s),\alpha)},$
where $J_{f,s}$ and $J_{f,u}$ are the Jacobian matrices of $f$ with respect to $s$ and $u$, respectively, and the overall NPV part is the component not included in the linearized nominal part, namely:
$\zeta(s,u_{NN}(s),\alpha)=f(s,Ks+u_{NN}(s),\alpha)-A_{K}s.$
@@
III-D Lyapunov-based LMI for robust stability
Using the QCs for the NN controller Lipschitz bound and system NPV sector bound, in [1], an LMI constraint is obtained to guarantee the existence of a robust Lyapunov function over a given state domain, Lipschitz constant, and parameter variation bound. The feasibility of this LMI constraint ensures the existence of a robust quadratic Lyapunov function to certify the the robust stability of the NN-controlled system and a RoS within the user-specified safety set that is invariant and ensures convergence to equilibrium, even under parameter variations.
@@ III-D Lyapunov-based LMI for robust stability
For the given system of (51), when the below LMI constraint of (61) is feasible for some $L,K,(P\in\mathbb{S}^{+}),(\Lambda\geq 0),(\gamma_{ij}\geq 0,\forall i\in 1,\ldots,m,j\in 1,\ldots,n)$, then there exists a robust quadratic Lyapunov function $\mathcal{V}(s)=s^{T}Ps$ such that $\pi(s)=Ks+\pi_{NN}(s)$ with $\pi_{NN}\in\Pi_{L}$ stabilizes the system of (51) over the given state domain under the variations of the parameters in its given domain:
@@
IV Robust, Optimal, Safe, Stabilizing Control for Quadcopter
IV-A Quadcopter dynamics & its Nominal control
@@
TABLE I: Quadcopter model parameters
@@
A quadcopter is an underactuated system where the degree of freedom for control is less than the number of states, and a typical control specification comprises a desired position ($x_{d},y_{d},z_{d}$) and a desired heading ($\psi_{d}$). We employ a cascade proportional-derivative (PD) controller as the nominal controller in our design. Using the position error, the positional P-controller computes the desired thrust ($T_{d}$) and desired roll/pitch angles ($\phi_{d},\theta_{d}$), which are then used to determine the attitude error which the attitude PD-controller uses to calculate the desired torque ($\tau_{d}$). The structure of this nominal controller is shown in Fig 2.
Figure 2: Nominal controller architecture
The control inputs are set to be the motor thrusts:
@@
The thrust and torque generated by the quadcopter’s rotors are proportional to the superposition of the square of rotor speeds. However, these values are influenced by aerodynamic conditions and external disturbances [22]. To account for these uncertainties, a parameter variation is introduced into the thrust and torque model that is then compensated through the added NN controller.
Table II represents the corresponding parameter variation bounds, indicating ±$\pm 5\%$ variation in the actual thrust and torque from their commanded values. Modeling these variations ensures that the model accounts for the real-world discrepancies arising from aerodynamic conditions, enhancing the robustness of the control design.
TABLE II: Quadcopter parameter variation bounds
@@
The system equation (101) consists of 12 states, 4 control inputs, and 4 parameter variations. The coordinates are shifted such that the equilibrium point is at the origin, i.e., $(x_{d},y_{d},z_{d},\psi_{d})=(0,0,0,0)$. With this coordinate shift, so that $(x_{d},y_{d},z_{d},\psi_{d})=(0,0,0,0)$, it follows that nominal Position and Attitude controllers are state-feedback.
Table III lists the parameter values for the nominal controller, which is chosen to ensure that the state matrix associated with the nominal controller, $A_{K}$, is stable with all its eigenvalues having negative real parts, i.e., the nominal controller stabilizes the linearized nominal system under zero parameter variation around the equilibrium point.
TABLE III: Nominal PD controller gains
@@
IV-B Maximal Lipschitz bound & Invariant safety domain
The feasibility of the LMI constraint (61) depends on system nonlinearity, safety domain, and parametric bound, and when feasible, one can find $(L,K,P,\Lambda,\{\lambda_{ij}\})$ that satisfy the LMI constraint so that the system (51) under the control $\pi(s)=Ks+\pi_{NN}(s);\,\pi_{NN}\in\Pi_{L}$ is certified as asymptotically stable with the Lyapunov function $\mathcal{V}=s^{T}Ps$ having a RS-RoS (Robust Safe Region of Stability) that is a sublevel set of $\mathcal{V}$. To maximize the class of robust, safe, and stable NN controllers, $L$ should be as large as possible, and to maximize the size of RS-RoS, the corresponding safety domain $\mathcal{S}$ should also be as large as possible. Maximal ($L,\mathcal{S}$) values are determined through the iterative search method of [1, Algorithm 1]. The process begins with an initial guess ($L_{0},\mathcal{S}_{0}$) set to small values, which for our quadcopter case is set as below:
$L_{0}=1.0;\mathcal{S}_{0}\!\equiv\!\left[\!\!\begin{array}[]{c}\mathcal{S}_{0}(x)=\mathcal{S}_{0}(y)=\mathcal{S}_{0}(z)\\ \mathcal{S}_{0}(\phi)=\mathcal{S}_{0}(\theta)=\mathcal{S}_{0}(\psi)\\ \mathcal{S}_{0}(\dot{x})=\mathcal{S}_{0}(\dot{y})=\mathcal{S}_{0}(\dot{z})\\ \mathcal{S}_{0}(p)=\mathcal{S}_{0}(q)=\mathcal{S}_{0}(r)\end{array}\!\!\right]\!=\!\left[\!\!\begin{array}[]{c}(-1.5,1.5)\\ (-0.2,0.2)\\ (-1.5,1.5)\\ (-0.2,0.2)\end{array}\!\!\right]\!.$
The ($L,\mathcal{S}$) values are iteratively increased, and the feasibility of the LMI constraint (61) is evaluated at each step. If the LMI constraint becomes infeasible, ($L,\mathcal{S}$) is reduced, and the feasibility is rechecked. This process is repeated with progressively smaller step sizes until the algorithm converges to the maximal Lipschitz bound and safety domain, denoted as ($L^{*},\mathcal{S}^{*}$), which in our case is found to be:
$L^{*}\!=\!1.2613;\mathcal{S}^{*}\!\!\equiv\!\!\!\left[\!\!\!\!\begin{array}[]{c}\mathcal{S}^{*}(x)\!=\!\mathcal{S}^{*}(y)=\!\mathcal{S}^{*}(z)\\ \mathcal{S}^{*}(\phi)\!=\!\mathcal{S}^{*}(\theta)\!=\!\mathcal{S}^{*}(\psi)\\ \mathcal{S}^{*}(\dot{x})\!=\!\mathcal{S}^{*}(\dot{y})=\!\mathcal{S}^{*}(\dot{z})\\ \mathcal{S}^{*}(p)\!=\!\mathcal{S}^{*}(q)\!=\!\mathcal{S}^{*}(r)\end{array}\!\!\!\!\right]\!\!\!=\!\!\!\left[\!\!\!\!\begin{array}[]{c}(-1.892,1.892)\\ (-0.252,0.252)\\ (-1.892,1.892)\\ (-0.266,0.266)\end{array}\!\!\!\!\right]\!\!.$
@@ IV-C Optimal NN controller training
To train an optimal NN controller satisfying the maximal Lipschitz bound $L^{*}$ found above so as to minimize reference tracking error and control cost, we employ Proximity Policy Optimization (PPO) [23], a deep reinforcement learning (RL) method based on an actor-critic structure. In an RL, an agent interacts with the environment to collect samples, where each sample is represented as $(s,u,r,\gamma,s^{\prime})$ where $s\in\mathcal{S},u\in\mathcal{U}$ are the current state and action ($\mathcal{S}$ and $\mathcal{U}$ are state and action domains), $r\in\mathbb{R}$ is the associated reward, $\gamma\in[0,1)$ is a discount factor, and $s^{\prime}\in\mathcal{S}$ is the next state. These samples represent the transition from $s$ to $s^{\prime}$ when the action $u$ is taken and the reward $r$ is received. Using the collected samples, policy and value, the neural networks used as function approximators, representing maps from state to action (actor network) and from state to value (value network) are trained, where value is the expected accumulated total rewards. (In contrast, the reward is an immediate return of a single action step.) The goal of RL is to optimize the network parameters to maximize the value.
Fig 3 illustrates this actor-critic framework used in our training of optimal NN controller. The environment models the quadcopter dynamics along with the nominal controller described in Equation (101). The NN controller provides control inputs that are fed into the environment (nominally controlled quadcopter), which executes the control and returns a “reward” value depending on the current state, and also advances the state. The “advantage” is defined as the difference between the state value and an action value, providing a measure of whether a particular action is better or worse than the default action suggested by the policy. We estimated the advantage using the method proposed in [24], and computed the policy gradient based on this estimated advantage.
Figure 3: Actor Critic Framework for RL training
The reward indicates how effectively the agent performs in achieving the desired objective, guiding the optimization of the control policy. In our case, the reward function minimizes the error between the reference state and the current state. When devising the control action, both the transient and steady-state responses are factored in: When the quadcopter is far from the reference position, its reward is computed from only the positional and directional errors. Once the quadcopter is close to the reference position, angular velocity, which must become small, is also factored in along with the positional and directional errors so as to reduce the oscillations around the reference. Given the position and direction errors defined as $p_{e}:=[x_{d}-x,y_{d}-y,z_{d}-z]^{T}$ and $\psi_{e}:=\psi_{d}-\psi$, and the angular velocity defined by $\nu:=[p,q,r]^{T}$, the reward function is formulated as:
@@
Figure 4: Actor and Critic networks for quadcopter control
Fig 4 shows the actor and critic network structures. In our design, the actor consists of five hidden layers, each containing 64 neurons with Rectified Linear Unit (ReLU) as the activation function. The final layer performs only linear operations without an activation function. The actor has two output branches: one computes the mean, and the other computes the variance of the action. Actions are then sampled from a multivariate normal distribution using the mean and variance during the training phase. During the testing phase, only the mean values are used to control the quadcopter. The critic network consists of four hidden layers, each with 192 or 256 neurons and ReLU activation function. This network evaluates the value function, guiding the optimization of the actor network.
At the beginning of each episode, the quadcopter’s position is initialized at the origin, and the reference position is randomly selected within a two-meter radius from the origin. Parameter variation ($\alpha$) is randomly sampled at every timestep from an uniform distribution within the given bounds, introducing uncertainties in thrust and torque. The simulation uses the first-order integration to update the state based on (101). Both actor and critic networks are trained using PPO. Table IV provides the PPO parameters we used for NN training. Definitions of these parameters can be found in [23], [24], and [25].
TABLE IV: PPO parameters used in Reinforcement Learning
@@
IV-D Certifying robust safety and stability
Robust safety is certified simply by starting within the maximal RS-ROS set $\mathcal{S}^{*}$ contained in the user-specified safety set, which by virtue of being a sublevel set of the computed robust Lyapunov function is invariant under system evolution, including random parametric variation, thereby preserving safety. Next, to certify the robust asymptotic stability of the closed-loop system, the Lipschitz constant of the NN controller must remain below the maximal computed Lipschitz bound $L^{*}$. It is therefore important to obtain a tight estimation of an NN’s Lipschitz constant. To achieve this, other than introducing L2-norm for Lipschitz continuity, we apply semi-definite programming based method [5] to estimate the Lipschitz constant of the trained actor network. In case the estimated Lipschitz bound of the trained NN is higher than the $L^{*}$ value, we simply scale down the weights of the final layer of the actor network (which by design is purely linear with no activation) so that the maximal Lipschitz bound requirement is satisfied. In our study, the Lipschitz constant of the trained actor network came out to be $L=3.1519>L^{*}=1.2613$. So we scaled down the weights of the final layer by a factor of $0.40020.4<L^{*}/L=$ 0.4002. After the corresponding scaling down, the Lipschitz bound of the NN controller is reduced to:
$L^{new}=1.2608<L^{*}=1.2613.$
Note since in our design, the final layer is a linear operation without an activation function, scaling it down does not alter the action pattern inferred by the actor network. Thereby this adjustment retains the NN controller’s functional behavior while also certifying its robust stability. The reduced conservativeness from the use of i) L2-norm for Lipschitz continuity, together with ii) semi-definite programming to estimate the Lipschitz bound of a NN, is demonstrated through the $L$ values in Table V, showing $>\!\!2400$-fold improvement.
TABLE V: Estimates of Lipschitz bound $L$
$L$ from $\infty$-norm [20] $L$ from L2-norm and using SDP[5]
2920 1.2608
V Evaluation of Synthesized Controller
We evaluate the performance of the controlled quadcopter under the nominal and NN controllers designed above. During the simulation of the controlled quadcopter, the $\alpha$ parameters are randomly drawn from the uniform distribution within the specified parameter bounds.